Consider the following PDF for a continus random variable f(x) X: 0 x<0,4 Calculate K Calculate P0, 1<x<0,3) Calculate P(X <= 0,2) Calculate E(X) Calculate Var(X) 3,75-Kx®2]
Answer #1

f(x) = 3.75-Kr for 0 < x < 0.4

For valid pdf,

0.4 f(x) dx1

0.4 (3.75 - Kr2) dr1

Rightarrow left [ 3.75x - Kx^3/3 ight ]^{0.4}_0 =1

3.75 * 0.4-K0.43/3 = 1

K(3.75 * 0.4- 1) *3/0.4

K 23.4375

0.3 P(0.1< X<0.3)3.75-23.43752) dr3.75 23.4375r/30.1 0.1

(3.75 * 0.3 -23.4375 0.33/3) (3.750.1 23.43750.13/3)

0.546875

P(X le 0.2) = int_{0}^{0.2}( 3.75 - 23.4375x^2 )~dx = left [ 3.75 x -23.4375x^3 / 3 ight ]^{0.2}_{0}

= (3.75 * 0.2 -23.4375 * 0.2^3 / 3)

= 0.6875

E(X) = int_{0}^{0.4} x f(x)~dx = int_{0}^{0.4}x( 3.75 - 23.4375x^2 )~dx

= int_{0}^{0.4}( 3.75x - 23.4375x^3 )~dx = left [3.75x^2/2 - 23.4375x^4/4 ight ]^{0.4}_0

3.75 0.42/2 23.4375 0.44/4-0.15

E(X^2) = int_{0}^{0.4} x^2 f(x)~dx = int_{0}^{0.4}x^2( 3.75 - 23.4375x^2 )~dx

0.4 0.4 -(3.75x2 - 23.4375r) dr3.75r/3-23.43755/50

3.75 0.4/3 - 23.4375 0.450.032

Var(X) = E(X^2) - (E(X))^2= 0.032 - 0.15^2 = 0.0095